I really need help soon!
Given f(x) = 2x^3 - 7 find f^-1(x)

Answer by stanbon(75887)

Given f(x) = 2x^3 - 7 find f^-1(x)
-------
1st: Interchange x and y to get::
x = 2y^3 - 7
-----
2nd: Solve for "y":
2y^3 = x + 7
---
y^3 = (x+7)/2
---
y = [(x+7)/2]^(1/3
-----------------------
That is f^-1(x)
=========================
Cheers,
Stan H.
-------------------------